OpenStudy (anonymous):

what is the square root of 45

OpenStudy (anonymous):

3root5

OpenStudy (saifoo.khan):

\[\sqrt {45} \to \sqrt{ 5 \times 9} \to \sqrt{ 5 \times 3 \times 3} \to 3 \sqrt 5\]

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):